OpenStudy (laddiusmaximus):

lim x->0 ((1=x)^3-1/x)

OpenStudy (anonymous):

1=x means 1+x?

OpenStudy (laddiusmaximus):

crap! yeah

OpenStudy (anonymous):

\[(1+x)^3-\frac{1}{x}\]so it looks like

OpenStudy (laddiusmaximus):

what?

OpenStudy (anonymous):

is that how the equation is written?

OpenStudy (laddiusmaximus):

no it is written (1+x)^3-1 / x\[\left( 1+x \right)^{3} -1\div x\]

OpenStudy (anonymous):

\[\frac{(1+x)^3-1}{x}\]so

OpenStudy (laddiusmaximus):

correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me type it out

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{(x+1-1)[(x+1)^2+(x+1)+1]}{x}\]

OpenStudy (anonymous):

That's because of a difference of 2 cubes

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{x(x^2+2x+1+x+2)}{x}=\lim_{x \rightarrow 0} x^2+3x+3\]

OpenStudy (anonymous):

plug in 0 for x from here, there's your solution

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

The limit is 3

OpenStudy (laddiusmaximus):

got you. Thank you!